# Noise-based reward-modulated learning

### v2

@@ Abstract
Biological neural systems efficiently learn from delayed rewards despite relying on noisy synaptic transmission and lacking centralized optimization mechanisms. In contrast, artificial neural networks trained with reinforcement learning typically rely on backpropagation (BP), which limits their use in resource-constrained systems or with non-differentiable components. While noise-based alternatives, like reward-modulated Hebbian learning (RMHL), provide a biologically grounded framework for credit assignment, they struggle with temporal delays and hierarchical processing —key challenges in real-world learning. In this work, we derive a novel noise-based learning rule to address these challenges. Drawing inspiration from biological neural circuits, our method uses reward prediction errors as its optimization target to generate increasingly advantageous behavior, and incorporates an eligibility trace to facilitate retrospective credit assignment. Its formulation relies on local information, aligning with biological constraints and enabling neuromorphic implementation. Experimental validation on reinforcement tasks (immediate and delayed rewards) shows our approach significantly outperforms RMHL and achieves performance comparable to BP, although with slower convergence due to its noise-driven updates. While tested on simple architectures, the results highlight the potential of noise-driven, brain-inspired learning for low-power adaptive systems, particularly in scenarios where energy efficiency and biological plausibility are a priority. These findings also offer mechanistic insights into how dopamine-like signals and synaptic stochasticity may jointly enable learning in biological networks, bridging computational models with neurobiological principles.
@@ 1 Introduction
The brain’s remarkable ability to learn and adapt to its environment is believed to arise from synaptic plasticity [4, 3, 37]. These local modifications allow neural circuits to acquire complex sensorimotor skills despite inherent challenges such as noisy neural communication [12], temporally delayed feedback/reward [23], and no centralized error-propagation mechanism to drive weight updates [48]. Understanding the mechanisms underlying this robust learning remains a central challenge in neuroscience and is increasingly inspiring novel approaches in artificial intelligence [11, 38].
When it comes to learning in artificial neural networks, backpropagation (BP) underpins modern deep learning [45]. However, its implementation in biological circuits is often seen as challenging [34]. BP requires precise, non-local error gradients, which contrasts with the locality of synaptic plasticity that depends on pre- and post-synaptic activity and neuromodulatory factors —akin to three-factor learning [21, 29]. Moreover, its deterministic nature is contrary to the inherent stochasticity of neural circuits [12, 51, 39], and its requirement for continuous, immediate error feedback contradicts biological learning, which often involves delayed feedback following a sequence of actions [33].
While some works have investigated how variants of backpropagation might be implemented in cortical microcircuits —employing predictive coding, feedback alignment, or dendritic error signals to carry nonlocal gradients [60, 59, 31]— others have proposed different plasticity rules inspired directly from neurobiology [46, 25, 63]. Among these approaches, noise‐based learning methods are especially promising. They leverage fluctuations in neural activity [19, 24, 61, 58, 16] or synaptic weights [13, 65, 7, 10], for estimating gradients. By injecting small random perturbations, these methods avoid the need for perfectly matched feedback pathways and rely only on information that biological synapses could plausibly access, aligning more closely with mechanisms observed in cortical neurons [50, 36].
@@
Given the prevalence of noise in biological neural systems [12, 51, 39], using noise as a mechanism for learning synaptic weights is an area of growing interest, with reward-modulated Hebbian learning (RMHL) [32, 40] as a promising candidate. RMHL, which is a form of three-factor Hebbian learning [21, 29], offers a potentially powerful mechanism for credit assignment without explicit backpropagation of errors. Despite theoretical advances, few noise-based methods have been adapted to real-world tasks. Some attempts which have been made include application to control problems [6], for spatiotemporal pattern generation [28], and for training non-differentiable spiking neural networks training [15], and these particularly struggle with delayed reward regimes. This gap has limited the integration of noise-based approaches for effective credit assignment in temporally extended setups.
@@
To bridge these gaps, we propose noise-based reward-modulated learning (NRL), a biologically-inspired, gradient-free learning method compatible with delayed feedback/rewards. NRL builds on established neuroscientific pillars. It employs eligibility traces [23, 26], which tag synapses based on recent pre/post activity and noise perturbations, enabling retrospective credit assignment over behavioral timescales. It also embraces reward prediction error (RPE), that is, dopamine-like signals [48, 47, 42] that modulate these traces to reinforce actions that yield unexpected positive rewards, driving learning toward increasingly more advantageous behaviors. We derive our learning rule from first principles, employing directional derivatives to compute a local gradient estimate at each synapse, rather than relying on explicit gradients. These directional derivatives are implemented through stochastic neurons, aligning with the inherently noisy nature of neural circuits and other physical systems. NRL estimates the directional derivative with two forward passes: a “noisy” pass with stochastic neurons and a “clean” pass without noise. In scenarios where a noiseless pass is infeasible, multiple noisy passes can be averaged to approximate the clean pass, maintaining performance.
We deliberately choose an abstract modelling framework so that our experiments isolate the learning rule’s behavior without introducing additional biological complexities. We employ rate-based neurons instead of spiking models, such that the dynamics reduce to smooth activation functions and continuous outputs. We also avoid distinguishing excitatory and inhibitory populations, allowing every unit to follow the same update equation. Finally, we treat potentiation and depression as two ends of a single weight-change continuum rather than separate processes, so that each synapse simply strengthens or weakens in proportion to its eligibility trace and reinforcement signal. In this way, we trade-off biological detail for more clear insight into capabilities and limitations of our proposed noise-based learning.
@@
We validate NRL on a suite of reinforcement-learning benchmarks spanning immediate-reward and delayed-reward tasks, demonstrating significantly superior learning efficiency when compared to RMHL (particularly when rewards were delayed by many steps), and competitive final performance compared to gradient-based baselines. Moreover, unlike RMHL, which breaks down in deeper architectures, NRL scales to multilayer networks, though converging more slowly due to its intrinsic stochasticity. Finally, we explore how these findings shed light on biological learning and point toward energy-efficient implementations on neuromorphic hardware, as well as novel, alternative, robust learning schemes for machine learning.
@@ 2.1 Neural system-environment interaction and reward signaling
We model a neural system within a dynamic environment, grounded in the reinforcement learning framework [54]. In this framework, the environment’s state at time $t$ is denoted by $s_{t}$. The neural system perceives $s_{t}$ and responds by initiating actions, $a_{t}$, which influence the environment. Each action or sequence of actions results in feedback in the form of positive or negative rewards $r_{t}$. This iterative process of perception, action, and outcome forms a feedback loop that enables the system to learn and adapt to the changing environment, ultimately seeking to maximize the rewards it receives.
The system’s decision-making is guided by a policy $\pi$, representing the probability of choosing an action, $a_{t}$, in a given state $s_{t}$ at time $t$, i.e., $\pi(a_{t}\mid s_{t})$. This policy captures the system’s learned strategy for choosing actions that are likely to yield favorable outcomes. To determine an action, the system computes a probability distribution across the set of possible actions, modelled as a categorical distribution $P(a_{t}\mid s_{t})\equiv\pi(a_{t}\mid s_{t})$ with $\sum_{a\in A}\pi(a\mid s_{t})=1$, where $A$ denotes the set of all possible actions.
The rewards $r_{t}$, obtained from the environment, serve as crucial learning signals. Internally, the system maintains a prediction of the expected reward, denoted by $\bar{r}_{t}$. While this prediction can be calculated through various mechanisms, we model it here for simplicity as a running average of recent rewards:
@@
This RPE signal is a central element of our model and draws a strong parallel to the phasic activity of dopamine neurons in the brain [48, 47, 42]. Dopamine neurons are known to exhibit increased activity for unexpected rewards and decreased activity for the unexpected absence of reward [17]. Although our running average is a simplified representation of how reward expectations are formed in biological systems, it effectively captures the fundamental dynamics of the RPE and its role as a learning signal. We will explore potential refinements of this reward prediction mechanism in the Discussion section.
@@ 2.2 Derivation of the learning rule
This section presents a complete derivation of the NRL update rule. We begin by establishing a gradient-based learning rule rooted in an optimization target that aims to maximize unexpected rewards from the environment by increasing the reward prediction error. Using RPE, rather than the direct reward, enables our system to adapt dynamically to changing environments by tracking how rewards deviate from expectations, improving long-term performance by continuously seeking rewards that exceed expectations, and maintaining robustness against variations in reward structures [54, 43]. Building on this foundation, we transition to a directional derivative framework, where stochastic noise within the network approximates gradients. This eliminates the need for explicit backpropagation and feedback phases, enabling learning through forward passes alone. Finally, we extend our noise-based approach to handle scenarios with delayed rewards, a hallmark of real-world problems. This extension enables NRL to adapt to environments where feedback is only obtained after a series of actions.
@@ 2.2.1 Gradient-based learning rule
The system is modelled as a multi-layered network of interconnected units, analogous to populations of neurons, which processes input representing sensory observations and produces a probability distribution over possible actions. For simplicity, we refer to these populations as ‘layers,’ where each layer performs a transformation of its input in two stages: first, a linear transformation, followed by a non-linear transformation. This can be expressed mathematically as
@@ 2.2.3 Learning with node level noise in neural network
Instead of directly perturbing the weights of the network $W^{l}_{t}$, we propose extending the learning rule by introducing noise directly into the neurons (nodes) of each layer. This strategy is advantageous because it can lead to reduced variance in gradient estimation, as perturbations occur in a lower-dimensional space (the neural activations) compared to the full weight space. This approach shares similarities with node perturbation (NP) [19, 24, 61, 58, 16], which has shown benefits in reducing gradient variance and offering more localized updates.
@@
$\tilde{x}^{l}_{t}=f\left(\tilde{h}^{l}_{t}+\xi^{l}_{t}\right)=f\left(W^{l}_{t}\tilde{x}^{l-1\top}_{t}+\xi^{l}_{t}\right)$ (9)
@@
$\nabla g\left(W^{l}_{t}\right)=\nabla g\left(h^{l}_{t}\right)\frac{\partial h^{l}_{t}}{\partial W^{l}_{t}}=\nabla g\left(h^{l}_{t}\right){{x}^{l-1\top}_{t}}.$ (11)
@@
$W^{l}_{t+1}\leftarrow W^{l}_{t}+\eta\delta_{t}\bar{\xi}^{l}_{t}{\rho}_{t}{\tilde{x}^{l-1\top}_{t}}$ (13)
@@ 2.2.4 Learning from sparse rewards
Let’s denote these discrete reward times as $\tau_{m}\in\{\tau_{0},\tau_{1},...\tau_{M}\}$. At these moments, we compute the RPE, $\delta_{\tau_{m}}$, as
@@
Here, $r_{\tau_{m}}$ is the reward received at time $\tau_{m}$, and $\bar{r}_{\tau_{m}}$ is the reward prediction, which is updated as the running average of recent rewards.
$\bar{r}_{\tau_{m}}=\bar{r}_{\tau_{m-1}}+\lambda\left(r_{\tau_{m}}-\bar{r}_{\tau_{m-1}}\right)$ (15)
@@ 2.3 Neural network architecture
$\tilde{x}^{l}_{t}$ $=f\left(W^{l}_{t}\tilde{x}^{l-1\top}_{t}+\xi^{l}_{t}\right)$ (18)
@@
$\tilde{y}_{t}$ $=s\left(W^{l}_{t}\tilde{x}^{l-1\top}_{t}+\xi^{l}_{t}\right)$ (19)
where $x^{l}_{t}$ represents the activity of units in layer $l$, $W^{l}_{t}$ is the matrix of synaptic weights connecting layer l1l-1 to layer $l$, and $\xi^{l}_{t}\sim\mathcal{N}(0,\sigma^{2}I_{m^{l}})$ represents random noise introduced into layer $l$, with $m^{l}$ being the number of units in such layer. This noisy propagation of activity is termed a “noisy pass.” The function $f(\cdot)$ is a non-linear activation function, which we implement as the LeakyReLU function such that $f(x)=x$ if $x\geq 0$ and $f(x)=\alpha x$ with $0<\alpha\ll 1$. The function $s(\cdot)$ represents a softmax transformation ${s}_{i}(x)=\ifrac{e^{x_{i}}}{\sum_{j=1}^{m^{l}}e^{x_{j}}}$, which transforms the network’s output into a probability distribution over possible actions. We also define a “clean pass,” representing the network’s activity in the absence of noise, by $x^{l}_{t}=f\left(W^{l}_{t}x^{l-1\top}_{t}\right)$ and $y_{t}=s\left(W^{l}_{t}x^{l-1\top}_{t}\right)$.
@@ 2.4 Experimental validation
$W^{l}_{\tau_{m}}\leftarrow W^{l}_{\tau_{m-1}}+\eta\left(\nabla_{W^{l}_{\tau_{m-1}}}\log\pi_{\theta_{\tau_{m-1}}}(a_{t}\mid s_{t})\right)\delta_{\tau_{m}}$ (20)
@@
Table 1: Comparison of the learning rules.
@@
NRL $W^{l}_{\tau_{m}}\leftarrow W^{l}_{\tau_{m-1}}+\eta\;\delta_{\tau_{m}}\!\sum_{t=\tau_{m-1}}^{\tau_{m}}\bar{\xi}^{\,l}_{t}\,\rho_{t}\,\tilde{x}^{\,l-1\top}_{t}$
BPTT $W^{l}_{\tau_{m}}\leftarrow W^{l}_{\tau_{m-1}}+\eta\left(\nabla_{W^{l}_{\tau_{m-1}}}\log\pi_{\theta_{\tau_{m-1}}}(a_{t}\mid s_{t})\right)\delta_{\tau_{m}}$
RMHL $W^{l}_{\tau_{m}}\leftarrow W^{l}_{\tau_{m-1}}+\eta\;\delta_{\tau_{m}}\!\sum_{t=\tau_{m-1}}^{\tau_{m}}\xi^{\,l}_{t}\,\tilde{x}^{\,l-1\top}_{t}$

### v3

@@ Abstract
The pursuit of energy-efficient and adaptive artificial intelligence (AI) has positioned neuromorphic computing as a promising alternative to conventional computing. However, achieving learning on these platforms requires techniques that prioritize local information while enabling effective credit assignment. Here, we propose noise-based reward-modulated learning (NRL), a novel synaptic plasticity rule that mathematically unifies reinforcement learning and gradient-based optimization with biologically-inspired local updates. NRL addresses the computational bottleneck of exact gradients by approximating them through stochastic neural activity, transforming the inherent noise of biological and neuromorphic substrates into a functional resource. Drawing inspiration from biological learning, our method uses reward prediction errors as its optimization target to generate increasingly advantageous behavior, and eligibility traces to facilitate retrospective credit assignment. Experimental validation on reinforcement tasks, featuring immediate and delayed rewards, shows that NRL achieves performance comparable to baselines optimized using backpropagation, although with slower convergence, while showing significantly superior performance and scalability in multi-layer networks compared to reward-modulated Hebbian learning (RMHL), the most prominent similar approach. While tested on simple architectures, the results highlight the potential of noise-driven, brain-inspired learning for low-power adaptive systems, particularly in computing substrates with locality constraints. NRL offers a theoretically grounded paradigm well-suited for the event-driven characteristics of next-generation neuromorphic AI.
@@ 1 Introduction
Modern artificial intelligence (AI) models are growing exponentially, leading to proportional increases in their computational demands and energy consumption [44]. This trend, combined with the growing requirement for widespread AI deployment in local edge devices [52, 60], creates a massive demand for systems that are both fast and energy-efficient. Meeting this demand requires a fundamental shift in computing architecture to overcome the energy and latency bottlenecks of conventional Von Neumann systems [59]. Neuromorphic computing [50, 45, 29] offers a promising path forward by mimicking the low energy, event-driven, and parallel processing capabilities of the brain [50, 35]. Although neuromorphic hardware is highly efficient for inference, the ability to perform effective on-chip learning and adaptation remains the primary challenge, lacking suitable on-chip learning algorithms.
Backpropagation (BP) remains the workhorse algorithm for training artificial neural networks, computing precise global gradients for weight updates [46]. However, its operational requirements are fundamentally incompatible with the constraints of neuromorphic hardware [10]. BP requires precise global error propagation, which translates into high-bandwidth global data movement and large memory requirements to store activations, the main source of energy consumption in deep learning systems [36]. This contrasts sharply with the hardware needs for local synaptic updates that rely only on information available at the synaptic weight. In addition, BP’s iterative forward and backward passes are incompatible with the asynchronous event-driven regime in which most neuromorphic devices operate [41], and its deterministic nature cannot tolerate inherit noise in neuromorphic circuits.
In response, local plasticity approaches and biologically-inspired learning rules have been proposed to bypass the constraints of BP [64, 63, 32, 47, 24, 67]. A particularly promising family of methods is noise-based learning, which leverages stochastic perturbations (either in weights [12, 69, 5, 9] or activations [18, 23, 65, 62, 15]) to approximate the true direction of the gradient. By injecting small random perturbations, these methods avoid the need for perfectly matched feedback pathways and can leverage inherit noise [51, 35].
@@
Given the prevalence of noise on neuromorphic and biological substrates [11, 53, 37], using noise as a mechanism to learn synaptic weights is an area of growing interest, with reward-modulated Hebbian learning (RMHL) [33, 38] as a promising candidate. RMHL, which is a form of three-factor Hebbian learning [20, 30], offers a potentially powerful mechanism for credit assignment without explicit backpropagation of errors. Despite theoretical advances, few noise-based methods have been adapted to real-world tasks. Some attempts which have been made include application to control problems [4], for spatiotemporal pattern generation [27], and for training non-differentiable spiking neural networks training [14], and these particularly struggle with delayed reward regimes. This gap has limited the integration of noise-based approaches for effective credit assignment in temporally extended setups.
@@
To bridge these gaps, we propose noise-based reward-modulated learning (NRL), a biologically-inspired, gradient-free learning method, which is compatible with delayed feedback/rewards. We derive our learning rule from first principles, employing directional derivatives to compute a local gradient estimate at each synapse. These directional derivatives are implemented through stochastic neurons, aligning with the noisy nature of neuromorphic physical systems, and are estimated with two forward passes: a “noisy” pass with stochastic neurons and a “clean” pass without noise. In scenarios where a noiseless pass is infeasible, multiple noisy passes can be averaged to approximate the clean pass, maintaining performance as shown in our experiments. NRL draws from established neuroscientific concepts, by employing eligibility traces [22, 25], which tag synapses based on recent pre/post activity and noise perturbations, allowing retrospective credit assignment over behavioral timescales. It also incorporates reward prediction errors (RPEs), a dopamine-like signal [49, 48, 40] that modulates eligibility traces to reinforce actions that yield unexpected positive rewards, driving learning toward increasingly more advantageous behaviors. We deliberately utilize a rate-based neural model, instead of a spiking model, to focus the analysis on the core computational capabilities and limitations of the learning rule itself, making our findings immediately applicable to neuromorphic substrates.
@@
We validate NRL on a suite of reinforcement-learning benchmarks that span immediate-reward and delayed-reward tasks, demonstrating significantly superior learning efficiency compared to RMHL (particularly when rewards were delayed by many steps) and competitive final performance compared to the gradient-based baseline. Moreover, unlike RMHL, which breaks down in deeper architectures, NRL scales to multilayer networks, though converging more slowly due to its intrinsic stochasticity.
Our findings argue that NRL represents a promising and scalable step toward developing low-energy and fast adaptive AI systems. By transforming the stochasticity of neuromorphic hardware from a challenge into a computational resource, NRL provides a crucial new algorithmic path for enabling complex, continuous learning at the edge, as well as novel, alternative, robust learning schemes for machine learning.
@@ 2.1 Neural system-environment interaction and reward signaling
We model a neural system within a dynamic environment, grounded in the reinforcement learning framework [56]. In this framework, the state of the environment at time $t$ is denoted by $s_{t}$. The neural system perceives $s_{t}$ and responds by initiating actions $a_{t}$, which influence the environment. Each action or sequence of actions results in feedback in the form of positive or negative rewards $r_{t}$. This iterative process of perception, action, and outcome forms a feedback loop that enables the system to learn and adapt to the changing environment, ultimately seeking to maximize the rewards it receives.
The decision-making of the system is guided by a policy $\pi$, which represents the probability of choosing an action, $a_{t}$, in a given state $s_{t}$ at time $t$, i.e., $\pi(a_{t}\mid s_{t})$. This policy captures the system’s learned strategy for choosing actions that are likely to yield favorable outcomes. To determine an action, the system computes a probability distribution across the set of possible actions, modeled as a categorical distribution $P(a_{t}\mid s_{t})\equiv\pi(a_{t}\mid s_{t})$ with $\sum_{a\in A}\pi(a\mid s_{t})=1$, where $A$ denotes the set of all possible actions.
The rewards $r_{t}$, obtained from the environment, serve as crucial learning signals. Internally, the system maintains a prediction of the expected reward, denoted by $\bar{r}_{t}$. Although this prediction can be calculated through various mechanisms, we model it here for simplicity as a running average of recent rewards as
@@
This RPE signal, which draws a strong parallel to the phasic activity of dopamine neurons in the brain [49, 48, 40, 16], functions as a low-bandwidth global feedback signal that is inherently compatible with the communication constraints of neuromorphic hardware. Although our running average is a simplified representation of how reward expectations are formed, it effectively captures the fundamental dynamics of the RPE and its role as an adaptive learning signal. We will explore potential refinements of this reward prediction mechanism in the Discussion section.
@@ 2.2 Derivation of the learning rule
This section presents a complete derivation of the NRL update rule. We begin by establishing a gradient-based learning rule rooted in an optimization target that aims to maximize unexpected rewards from the environment by increasing the reward prediction error. Using RPE, rather than direct reward, enables our system to adapt dynamically to changing environments by tracking how rewards deviate from expectations, improving long-term performance by continuously seeking rewards that exceed expectations, and maintaining robustness against variations in reward structures [56, 42].
Building on this foundation, we transition to a directional derivative framework, where intrinsic noise within the network is leveraged to approximate gradients. This critical step eliminates the need for backpropagation and feedback phases, enabling learning through forward passes alone. Finally, we extend our noise-based approach to handle scenarios with delayed rewards, a hallmark of real-world problems. This extension enables NRL to adapt to environments where feedback is obtained only after a series of actions.
@@ 2.2.1 Gradient-based learning rule
The system is modeled as a multi-layered network of interconnected units, analogous to populations of neurons, which processes input representing sensory observations and produces a probability distribution over possible actions. For simplicity, we refer to these populations as ‘layers,’ where each layer performs a transformation of its input in two stages: first, a linear transformation, followed by a non-linear transformation. This can be expressed mathematically as
@@ 2.2.3 Learning with node level noise in neural network
Instead of directly perturbing the weights of the network $W^{l}_{t}$, we propose introducing noise directly into the neurons (nodes) of each layer. This strategy is advantageous because it leads to reduced variance in gradient estimation, as perturbations occur in a lower-dimensional space (the neural activations) compared to the full weight space. In addition, it naturally reduces the computational cost and communication overhead associated with high-dimensional perturbations. This approach shares similarities with node perturbation (NP) [18, 23, 65, 62, 15], which has shown benefits in reducing gradient variance and offering more localized updates.
@@
$\tilde{x}^{l}_{t}=f\left(\tilde{h}^{l}_{t}+\xi^{l}_{t}\right)=f\left(W^{l}_{t}(\tilde{x}^{l-1}_{t})^{\top}+\xi^{l}_{t}\right)$ (9)
@@
$\nabla g\left(W^{l}_{t}\right)=\nabla g\left(h^{l}_{t}\right)\frac{\partial h^{l}_{t}}{\partial W^{l}_{t}}=\nabla g\left(h^{l}_{t}\right)(x^{l-1}_{t})^{\top}.$ (11)
@@
$W^{l}_{t+1}\leftarrow W^{l}_{t}+\eta\delta_{t}\bar{\xi}^{l}_{t}{\rho}_{t}(\tilde{x}^{l-1}_{t})^{\top}$ (13)
@@ 2.2.4 Learning from sparse rewards
Let us denote these discrete reward times as $\tau_{m}\in\{\tau_{0},\tau_{1},...\tau_{M}\}$. At these moments, we compute the RPE as
@@
Here, $r_{\tau_{m}}$ is the reward received at time $\tau_{m}$, and $\bar{r}_{\tau_{m}}$ is the reward prediction, which is updated as the running average of recent rewards:
$\bar{r}_{\tau_{m}}=\bar{r}_{\tau_{m-1}}+\lambda\left(r_{\tau_{m}}-\bar{r}_{\tau_{m-1}}\right).$ (15)
@@ 2.3 Neural network architecture
$\tilde{x}^{l}_{t}$ $=f\left(W^{l}_{t}(\tilde{x}^{l-1}_{t})^{\top}+\xi^{l}_{t}\right)$ (18)
@@
$\tilde{y}_{t}$ $=s\left(W^{l}_{t}(\tilde{x}^{l-1}_{t})^{\top}+\xi^{l}_{t}\right)$ (19)
where $x^{l}_{t}$ represents the activity of units in layer $l$, $W^{l}_{t}$ is the matrix of synaptic weights connecting layer l1l-1 to layer $l$, and $\xi^{l}_{t}\sim\mathcal{N}(0,\sigma^{2}I_{m^{l}})$ represents random noise introduced into layer $l$, with $m^{l}$ being the number of units in such layer. This noisy propagation of activity is termed a “noisy pass.” The function $f(\cdot)$ is a non-linear activation function, which we implement as the LeakyReLU function such that $f(x)=x$ if $x\geq 0$ and $f(x)=\alpha x$ with $0<\alpha\ll 1$. The function $s(\cdot)$ represents a softmax transformation ${s}_{i}(x)=\ifrac{e^{x_{i}}}{\sum_{j=1}^{m^{l}}e^{x_{j}}}$, which transforms the network’s output into a probability distribution over possible actions. We also define a “clean pass,” representing the network’s activity in the absence of noise, by $x^{l}_{t}=f\left(W^{l}_{t}(x^{l-1}_{t})^{\top}\right)$ and $y_{t}=s\left(W^{l}_{t}(x^{l-1}_{t})^{\top}\right)$.
@@ 2.4 Experimental validation
$W^{l}_{\tau_{m}}\leftarrow W^{l}_{\tau_{m-1}}+\eta\left(\nabla_{W^{l}_{\tau_{m-1}}}\log\pi_{\theta_{\tau_{m-1}}}(a_{t}\mid s_{t})\right)\delta_{\tau_{m}}.$ (20)
@@
Table 1: Learning rules. The terms in the BPTT update are rearranged for easier comparison.
@@
NRL $W^{l}_{\tau_{m}}\leftarrow W^{l}_{\tau_{m-1}}+\eta\;\delta_{\tau_{m}}\!\sum_{t=\tau_{m-1}}^{\tau_{m}}\bar{\xi}^{\,l}_{t}\,\rho_{t}\,(\tilde{x}^{l-1}_{t})^{\top}$
BPTT $W^{l}_{\tau_{m}}\leftarrow W^{l}_{\tau_{m-1}}+\eta\delta_{\tau_{m}}\nabla_{W^{l}_{\tau_{m-1}}}\log\pi_{\theta_{\tau_{m-1}}}(a_{t}\mid s_{t})$
RMHL $W^{l}_{\tau_{m}}\leftarrow W^{l}_{\tau_{m-1}}+\eta\;\delta_{\tau_{m}}\!\sum_{t=\tau_{m-1}}^{\tau_{m}}\xi^{\,l}_{t}\,(\tilde{x}^{l-1}_{t})^{\top}$